OpenStudy (anonymous):

find three consecutive numbers whose sum is 99

OpenStudy (anonymous):

32+33+34=99

OpenStudy (anonymous):

i need work shown step by step

OpenStudy (anonymous):

N + (N+1) + (N+2) = 99 3N + 3 = 99 3N = 96 N = 32 32, 33, 34